)Jim Smith owns a rental property which had a UCC of $85,000 at the beginning of 20x0. After all allowable expenses other than CCA, Jim's total rental income was $1,000 in 20x0 and $10,000 in 20x1. Jim always deducts the maximum CCA allowed. What is the UCC for his rental property at the end of 20x1? (The property is a Class 1 building amortized at 4%.)
Answer #1

Answer: 80,640

Calculation

20X0
UCC at beginning of 20X0 85,000
CCA -1,000
UCC at end of 20X0 84,000
20X1
UCC at beginning of 20X1 84,000
CCA -3,360
UCC 80,640

CCA calculation :

20X0 = 85000*0.04*0.5 = 1,700, which is restricted to $1000 rental income

20X1 = 84,000*0.04 = 3,360 which is less than rental income of $10,000.